OpenStudy (llroberts43):

I do not understand how to graph 3x-y<4. I thought I was supposed to substitute 0 for x and solve for y and then substitute 0 for y and solve for x to come up with the intercepts. When I do that I end up with y=4 (0,4) and then I end up with 3x = 4 and I don't know what to do next because 3 won't go into 4.

jimthompson5910 (jim_thompson5910):

to graph 3x-y<4, we'll need to first graph 3x-y=4 which is the boundary line

jimthompson5910 (jim_thompson5910):

plugging in x = 0 leads to 3x-y=4 3*0-y = 4 0-y = 4 -y = 4 y = -4 so it should be (0,-4) instead of (0,4)

OpenStudy (llroberts43):

Oh, I didn't know what to do with the negative sign so i ignored it

jimthompson5910 (jim_thompson5910):

now let's plug in another x value, say x = 1 3x-y=4 3*1-y = 4 3-y = 4 solve for y to get y = ??

OpenStudy (llroberts43):

What do I do with the 3?

jimthompson5910 (jim_thompson5910):

to move that over, you subtract 3 from both sides

OpenStudy (llroberts43):

-1

jimthompson5910 (jim_thompson5910):

yes if x = 1, then y = -1

jimthompson5910 (jim_thompson5910):

so the two points (0,-4) and (1,-1) are on the graph of 3x-y=4

jimthompson5910 (jim_thompson5910):

plot the two points and draw a straight line through both of them. Extend the line out as far as you can in both directions

OpenStudy (llroberts43):

thank you. It's always something so small that I do wrong.

jimthompson5910 (jim_thompson5910):

do you know how do to the shaded region?

OpenStudy (llroberts43):

yes

jimthompson5910 (jim_thompson5910):

ok great, just checking